# A Novel Multiscale Spatial Phase Estimate with the Structure Multivector

### v1

@@ Abstract
The monogenic signal (MS) was introduced by Felsberg and Sommer [3], and independently by Larkin [7] under the name vortex operator. It is a two-dimensional (2D) analog of the well-known analytic signal, and allows for direct amplitude and phase demodulation of (amplitude and phase) modulated images so long as the signal is intrinsically one-dimensional (i1D). Felsberg’s PhD dissertation also introduced the structure multivector (SMV), a model allowing for intrinsically 2D (i2D) structure. While the monogenic signal has become a well-known tool in the image processing community, the SMV is little used, although even in the case of i1D signals it provides a more robust orientation estimation than the MS. We argue the SMV is more suitable in standard i1D image feature extraction due to the this improvement, and extend the steerable wavelet frames of Held et al. [4] to accommodate the additional features of the SMV. We then propose a novel quality map based on local orientation variance that yields a multiscale phase estimate which performs well even when SNR $\geq 1$. The performance is evaluated on several synthetic phase estimation tasks as well as on a fine-scale fingerprint registration task related to the 2D phase demodulation problem.
@@ 1 Introduction
Many problems in imaging science rely on spatial phase measurements, e.g. 2D interferometry, interferometric SAR (InSAR), and require a preprocessing step to estimate the true spatial phase of an image or set of images [12]. Any improvement to this estimate will thus improve downstream analysis. A standard approach for estimating spatial phase of images is to use the phase of the monogenic signal [6] [13]. The first improvement to this estimate is to produce a multiscale monogenic phase estimate, see Kaseb et al. [5], for instance, which makes use of isotropic wavelets [4] [10], and provides a robust phase estimate in the presence of image corruption. The main contributions of this article are: 1) to employ the structure multivector (SMV) in place of the monogenic signal in order to extract a more robust feature set at any given scale; and 2) to define a novel quality measure at each scale, based on the features of the SMV, in order to determine the optimal local feature set around a given point in an image. We perform several experiments on synthetic images to showcase the application of the our multiscale phase estimation, and further solve phase and amplitude demodulation problem in 2D to display the utility of this estimate. Additionally, we use our multiscale phase estimate in order to solve a fine-scale fingerprint registration problem as described in [1]. Lastly, we have provided a Julia module that includes the code needed to reproduce any figures and experiments shown in this paper, as well as standalone functions to perform our multiscale phase estimation: https://gitlab.com/briancknight/SSVM2025.
The rest of the article is organized as follows. In Section 2 we discuss the standard signal model used to describe fringe patterns, and define the monogenic signal (MS), steerable wavelets, and the structure multivector (SMV). Section 3 describes the novel multiscale phase estimate derived from the SMV features at each scale, and Section 4 concludes with numerical experiments showcasing the improved phase estimation, as well as improved accuracy in 2D phase and amplitude demodulation tasks. Section 5 provides a brief conclusion to the article. In the additional appendix, Appendix 0.A provides more details in the construction of the structure multivector, and Appendix 0.B provides an error analysis for the feature set of the SMV when the i2D signal model is violated.
@@ 2.1 The i1D Signal Model and the Monogenic Signal
The monogenic signal is a quaternion valued signal that extends the real-valued 2D signal via the Riesz transform, the appropriate 2D extension of the Hilbert transform. If our original signal obeys the signal model
@@
$\begin{split}f_{M}({\boldsymbol{x}})&=f({\boldsymbol{x}})+\mathcal{R}f({\boldsymbol{x}})\\ &=f({\boldsymbol{x}})+{\boldsymbol{i}}\mathcal{R}_{1}f({\boldsymbol{x}})+{\boldsymbol{j}}\mathcal{R}_{2}f({\boldsymbol{x}}),\end{split}$
@@
is the Riesz transform for $x_{k}$ direction in ${\mathbb{R}}^{2}$. Larkin [7] showed that for signals of the i1D model described above, the Riesz transform obeys the asymptotic Bedrosian Principle:
@@ 2.3 The i2D Signal Model and the Structure Multivector
Two-dimensional signals of course can vary in two or more orientations in any given local patch, and further work has been done using hypercomplex signal processing that can deal with these cases. One extension, the structure multivector (SMV), was introduced along with the monogenic signal in the PhD dissertation of Felsberg [3]. It is designed to deal with signals of the form
@@
The features of the SMV will essentially be that of two monogenic signals, and to accommodate these additional features the SMV lives in a larger dimensional Clifford algebra, $\text{C}\ell_{3,0}$ which subsumes the quaternions. This algebra is generated by the orthonormal basis $e_{1},e_{2},e_{3}$ satisfying the relations $e_{i}e_{j}+e_{j}e_{i}=2\delta_{ij}$. In general we denote the product $e_{i_{1}}e_{i_{2}}\cdots e_{i_{n}}:=e_{i_{1}i_{2}\cdots i_{n}}$. For more information on Clifford algebras and the construction of the SMV see [2], [3] and Appendix 0.A. We only give here the minimum details needed to construct the SMV. We consider an image of the form $\mathbf{f}:{\mathbb{R}}^{2}\to e_{3}{\mathbb{R}}$, $\mathbf{f}({\boldsymbol{x}})=f(x,y)e_{3}.$
@@
$\begin{split}M_{S}({\boldsymbol{x}})&=\left[\mathbf{f}({\boldsymbol{x}})+(h_{2}^{1}*\mathbf{f})({\boldsymbol{x}})\right]+e_{3}\left[(h_{2}^{2}*\mathbf{f})({\boldsymbol{x}})+(h_{2}^{3}*\mathbf{f})({\boldsymbol{x}})\right]\\ &=M_{0}+M_{1}e_{1}+M_{2}e_{2}+M_{3}e_{3}+M_{23}e_{23}+M_{31}e_{31}+M_{12}e_{12}\end{split}$ (3)
@@ 3.1 A Multiscale Phase Estimate Using the Structure Multivector
In [5] $\mathcal{Q}^{(k)}({\boldsymbol{x}})$ is defined to be the amplitude of the monogenic signal at that scale. The analog in this paper is the major amplitude of the SMV at each scale, which we call the local amplitude quality. The idea is to choose the scale with the maximum local energy; if the underlying signal in question is some sort of well structured fringe pattern, such as an interferogram or fingerprint, this scale should correspond to the true spatial phase to be estimated and the multiscale feature set should be robust to noise or local signal corruption when the SNR is sufficiently well behaved.
When the local SNR is close to 1, however, the scale with the dominant amplitude is likely to be the noise itself. Still, it may be assumed that the signal corruption does not contain coherent structural information, and so we submit that a local quality metric which makes use of the local structural information should enable good phase estimation even when SNR $\geq 1$.
@@ 4.1 Multiscale Phase Estimation Experiments
Given a noisy plane wave of the form $f({\boldsymbol{x}})=\cos(\omega({\boldsymbol{n}}\cdot{\boldsymbol{x}}))+\eta_{\sigma}({\boldsymbol{x}})$, where $\eta_{\sigma}({\boldsymbol{x}})$ is a Gaussian random variable with mean zero and standard deviation $\sigma$, the goal is to estimate the true phase function $\omega({\boldsymbol{n}}\cdot{\boldsymbol{x}})\mod 2\pi$. In our experiment we discretize so that ${\boldsymbol{x}}[i,j]=[-\pi+\frac{2\pi i}{N},-\pi+\frac{2\pi j}{N}]$, for $i,j=0,N-1$, for $N=2^{M}$, $\omega$ ranges between $2^{3}$ and $2^{M-2}$, and ${\boldsymbol{n}}=[\cos(\pi/4),\sin(\pi/4)]^{T}$. We use the structural similary index measure (SSIM) [11] to compare the quality of the estimated phase to the ground truth. Figure 5 shows the quality of the estimated phase for $0\leq\sigma\leq 1.5$. Because the true amplitude of the noiseless signal is 1 everywhere, $\sigma$ can be thought of as the SNR, hence $\sigma=1$ is the point at which the Gaussian noise begins to dominate the underlying plane wave structure. The amplitude-based multiscale phase of the monogenic signal and SMV perform well until $\sigma$ surpasses 0.75, beyond which the estimate is unusable. In contrast, the estimate from the local orientation variance quality and the estimate given by the product of the amplitude and local orientation variance quality provide high-quality phase estimates well beyond this point.
Figure 4: Example of a (a) noisy plane wave and (b) the true underlying phase, and (c) a parabolic chirp signal along with (d) it’s true phase
@@ 4.2 Phase Demodulation with Multiscale Major Phase
The multiscale major phase estimate given by the the multiscale SMV representation of a signal $f$ motivates further experiments in phase demodulation. In two dimensions, the phase demodulation problems can be stated as follows: Given the carrier wave $c({\boldsymbol{x}})=A\cos(\omega_{c}{\boldsymbol{n}}\cdot{\boldsymbol{x}}+\phi_{c})$, and a message $m({\boldsymbol{x}})$ where $\hat{m}({\boldsymbol{u}})=0$ for $|{\boldsymbol{u}}|\geq\omega_{c}$, the phase modulated (PM) signal is given by:
@@
Typically the message is also sinusoidal in structure. For a baseline phase demodulation task, we attempt to recover a sinusoidal message after phase modulation and noise corruption. In this case, motivated by the phase estimation of the parabolic chirp, we use the overcomplete feature set when computing the multiscale phase to more accurately capture the overlapping frequency bands resulting from the phase modulation. Figure 8 shows the noisy phase modulated message, the ground truth message, and recovered messages via an amplitude based mulstiscale phase estimate versus the proposed product method. Again our tests conclude the product quality map allows for accurate phase estimation even when SNR is large.
@@ 4.3 Application to Deformable Fingerprint Registration
Global phase unwrapping algorithms typically have several major discontinuities which provide unreliable frequency information. In our experiments we compute a local frequency estimate $\omega({\boldsymbol{x}})$ by differentiation of a windowed phase unwrapping of $\Delta\phi({\boldsymbol{x}})$. The phase and orientation values used are those provided by our multiscale method. Figure 10 gives an example of the fine-scale registration produced by this method.
@@
Table 1 expresses the quality of the fingerprint match after successful rigid registration of fingerprints (from FVC2004 DB1-B [8]), and then after the additional fine-scale deformable registration is applied. Because we are interested only in the performance of the fine-scale registration algorithm, we add noise only after successful rigid registration. We find that even at large noise levels we are able to estimate accurate fine-scale registration in well structured areas of the fingerprints
@@ Appendix 0.A Construction of the SMV
Two-dimensional signals of course can vary in two or more orientations in any given local patch, and further work has been done using hypercomplex signal processing that can deal with these cases. One extension, the structure multivector (SMV), was introduced along with the monogenic signal in the PhD dissertation of Felsberg [3]. It is designed to deal with signals of the form
@@
The features of the SMV will essentially be that of two monogenic signals, and to accommodate these additional features the SMV lives in a larger dimensional Clifford algebra, $\text{C}\ell_{3,0}$ which subsumes the quaternions. For more information on Clifford algebras see [2], [3]. This algebra is generated by the orthonormal basis $e_{1},e_{2},e_{3}$ satisfying the relations $e_{i}e_{j}+e_{j}e_{i}=2\delta_{ij}$. In general we denote the product $e_{i_{1}}e_{i_{2}}\cdots e_{i_{n}}:=e_{i_{1}i_{2}\cdots i_{n}}$.

### v2

# A Structurally Coherent Spatial Phase Estimate

@@ Abstract
The monogenic signal (MS) was introduced by Felsberg and Sommer [3], and independently by Larkin [7] under the name vortex operator. It is a two-dimensional (2D) analog of the well-known analytic signal, and allows for direct amplitude and phase demodulation of (amplitude and phase) modulated images so long as the signal is intrinsically one-dimensional (i1D). Felsberg’s PhD dissertation also introduced the structure multivector (SMV), a model allowing for intrinsically 2D (i2D) structure. While the monogenic signal has become a well-known tool in the image processing community, the SMV is little used, although even in the case of i1D signals it provides a more robust orientation estimation than the MS. We argue the SMV is more suitable in standard i1D image feature extraction due to the this improvement, and extend the steerable wavelet frames of Held et al. [4] to accommodate the additional features of the SMV. We then propose a novel quality map based on local orientation variance which values structurally coherent patches. This yields a multiscale phase estimate which performs well even when signal to noise ratio (SNR) is $\leq 1$. The performance is evaluated on several synthetic phase estimation tasks as well as on a fine-scale fingerprint registration task related to the 2D phase demodulation problem.
@@ 1 Introduction
Many problems in imaging science rely on spatial phase measurements, e.g. 2D interferometry, interferometric SAR (InSAR), and require a preprocessing step to estimate the true spatial phase of an image or set of images [12]. Any improvement to this estimate will thus improve downstream analysis. A standard approach for estimating spatial phase of images is to use the phase of the monogenic signal [6] [13]. The first improvement to this estimate is to produce a multiscale monogenic phase estimate, see Kaseb et al. [5], for instance, which makes use of isotropic wavelets [4] [10], and provides a robust phase estimate in the presence of image corruption. The main contributions of this article are: 1) to employ the structure multivector (SMV) in place of the monogenic signal in order to extract a more robust feature set at any given scale; and 2) to define a novel quality measure at each scale, based on the features of the SMV, in order to determine the optimal local feature set around a given point in an image. We perform several experiments on synthetic images to showcase the application of the our multiscale phase estimation, and further solve a phase and amplitude demodulation problem in 2D to display the utility of this estimate. Additionally, we use our multiscale phase estimate in order to solve a fine-scale fingerprint registration problem as described in [1]. Lastly, we have provided a Julia module that includes the code needed to reproduce any figures and experiments shown in this paper, as well as standalone functions to perform our multiscale phase estimation: https://gitlab.com/briancknight/SSVM2025.
The rest of the article is organized as follows. In Section 2 we discuss the standard signal model used to describe fringe patterns, and define the monogenic signal (MS), steerable wavelets, and the structure multivector (SMV). Section 3 describes the novel multiscale phase estimate derived from the SMV features at each scale, and Section 4 concludes with numerical experiments showcasing the improved phase estimation as well as improved accuracy in 2D phase and amplitude demodulation tasks, including one used in fine-scale fingerprint registration. Section 5 provides a brief conclusion to the article. In the additional appendix, Appendix 0.A provides more details in the construction of the structure multivector, and Appendix 0.B provides an error analysis for the feature set of the SMV when the i2D signal model is violated.
@@ 2.1 The i1D Signal Model and the Monogenic Signal
In 1D signal processing, a real-valued signal can be extended to a complex valued signal via the Hilbert transform, and this complex-valued extension can provide useful insight into the signal’s local amplitude and frequency. In 2D, the monogenic signal is a quaternion valued signal that extends the real-valued 2D signal via the Riesz transform,the appropriate 2D extension of the Hilbert transform. If our original signal obeys the signal model
@@
$\begin{split}f_{M}({\boldsymbol{x}})&=f({\boldsymbol{x}})+{\boldsymbol{i}}\mathcal{R}_{1}f({\boldsymbol{x}})+{\boldsymbol{j}}\mathcal{R}_{2}f({\boldsymbol{x}}),\end{split}$
@@
is the Riesz transform for the $x_{k}$ direction in ${\mathbb{R}}^{2}$, and $\mathcal{R}f({\boldsymbol{x}})={\boldsymbol{i}}\mathcal{R}_{1}f({\boldsymbol{x}})+{\boldsymbol{j}}\mathcal{R}_{2}f({\boldsymbol{x}})$ is the (total) Riezs transform. Larkin [7] showed that for signals of the i1D model described above, the Riesz transform obeys the asymptotic Bedrosian Principle:
@@ 2.3 The i2D Signal Model and the Structure Multivector
Two-dimensional signals can, of course, vary in two or more orientations in any given local patch, and further work has been done using hypercomplex signal processing that can deal with these cases. One extension, the structure multivector (SMV), was introduced along with the monogenic signal in the PhD dissertation of Felsberg [3]. It is designed to deal with signals of the form
@@
The features of the SMV will essentially be that of two monogenic signals, and to accommodate these additional features the SMV lives in a larger dimensional Clifford algebra, $\text{C}\ell_{3,0}$ which subsumes the quaternions. This algebra is generated by the orthonormal basis $e_{1},e_{2},e_{3}$ satisfying the relations $e_{i}e_{j}+e_{j}e_{i}=2\delta_{ij}$, and consists of $82^{3}=$ 8 elements: $1,e_{1},e_{2},e_{3},e_{12},e_{23},e_{31},e_{123}$. In general we denote the product $e_{i_{1}}e_{i_{2}}\cdots e_{i_{n}}:=e_{i_{1}i_{2}\cdots i_{n}}$. For more information on Clifford algebras and the construction of the SMV see [2], [3] and Appendix 0.A. We only give here the minimum details needed to construct the SMV. We consider an image of the form $\mathbf{f}:{\mathbb{R}}^{2}\to e_{3}{\mathbb{R}}$, $\mathbf{f}({\boldsymbol{x}})=f(x,y)e_{3}.$
@@
$\begin{split}M_{S}({\boldsymbol{x}})&=\left[\mathbf{f}({\boldsymbol{x}})+(h_{2}^{1}*\mathbf{f})({\boldsymbol{x}})\right]+e_{3}\left[(h_{2}^{2}*\mathbf{f})({\boldsymbol{x}})+(h_{2}^{3}*\mathbf{f})({\boldsymbol{x}})\right]\\ &=M_{0}+M_{1}e_{1}+M_{2}e_{2}+M_{3}e_{3}+M_{23}e_{23}+M_{31}e_{31}+M_{12}e_{12}.\end{split}$ (3)
The explicit definitions of these functions are given below, where ${\boldsymbol{x}}=xe_{1}+ye_{2}$:
$M_{1}$ $=\frac{x}{2\pi|{\boldsymbol{x}}|^{3}}*f({\boldsymbol{x}}),$ $M_{2}$ $=\frac{y}{2\pi|{\boldsymbol{x}}|^{3}}*f({\boldsymbol{x}}),$ $M_{3}$ $=f({\boldsymbol{x}}),$
$M_{23}$ $=\frac{3(3x^{2}y-y^{3})}{2\pi|{\boldsymbol{x}}|^{5}}*f({\boldsymbol{x}}),$ $M_{31}$ $=\frac{3(3xy^{2}-x^{3})}{2\pi|{\boldsymbol{x}}|^{5}}*f({\boldsymbol{x}}),$
$M_{0}$ $=\frac{-2(x^{2}-y^{2})}{2\pi|{\boldsymbol{x}}|^{4}}*f({\boldsymbol{x}}),$ $M_{12}$ $=\frac{-4xy}{2\pi|{\boldsymbol{x}}|^{4}}*f({\boldsymbol{x}}).$
@@ 3.1 A Multiscale Phase Estimate Using the Structure Multivector
In [5] $\mathcal{Q}^{(k)}({\boldsymbol{x}})$ is defined to be the amplitude of the monogenic signal at that scale. The analog in this paper is the major amplitude of the SMV at each scale, which we call the local amplitude quality. The idea is to choose the scale with the maximum local energy; if the underlying signal in question is some sort of well structured fringe pattern, such as an interferogram or fingerprint, this scale should correspond to the true spatial phase to be estimated and the multiscale feature set should be robust to noise or local signal corruption when the signal to noise ratio (SNR) is sufficiently well behaved.
When the local SNR is close to 1, however, the scale with the dominant amplitude is likely to be the noise itself. Still, it may be assumed that the signal corruption does not contain coherent structural information, and so we submit that a local quality metric which makes use of the local structural information should enable good phase estimation even when SNR $\leq 1$.
@@ 4.1 Multiscale Phase Estimation Experiments
Given a noisy plane wave of the form $f({\boldsymbol{x}})=\cos(\omega({\boldsymbol{n}}\cdot{\boldsymbol{x}}))+\eta_{\sigma}({\boldsymbol{x}})$, where $\eta_{\sigma}({\boldsymbol{x}})$ is a Gaussian random variable with mean zero and standard deviation $\sigma$, the goal is to estimate the true phase function $\omega({\boldsymbol{n}}\cdot{\boldsymbol{x}})\mod 2\pi$. In our experiment we discretize so that ${\boldsymbol{x}}[i,j]=[-\pi+\frac{2\pi i}{N},-\pi+\frac{2\pi j}{N}]$, for $i,j=0,N-1$, for $N=2^{M}$, $\omega$ ranges between $2^{3}$ and $2^{M-2}$, and ${\boldsymbol{n}}=[\cos(\pi/4),\sin(\pi/4)]^{T}$. We use the structural similary index measure (SSIM) [11] to compare the quality of the estimated phase to the ground truth. Figure 5 shows the quality of the estimated phase for $0\leq\sigma\leq 1.5$. Because the true amplitude of the noiseless signal is 1 everywhere, $\sigma$ can be thought of as the recipricol of SNR, hence $\sigma=1$ is the point at which the Gaussian noise begins to dominate the underlying plane wave structure. The amplitude-based multiscale phase of the monogenic signal and SMV perform well until $\sigma$ surpasses 0.75, beyond which the estimate is unusable. In contrast, the estimate from the local orientation variance quality and the estimate given by the product of the amplitude and local orientation variance quality provide high-quality phase estimates well beyond this point.
Figure 4: Example of a (a) noisy plane wave and (b) the true underlying phase, and (c) a parabolic chirp signal along with (d) its true phase
@@ 4.2 Phase Demodulation with Multiscale Major Phase
The multiscale major phase estimate given by the the multiscale SMV representation of a signal $f$ motivates further experiments in phase demodulation. In two dimensions, the phase demodulation problem can be stated as follows: Given the carrier wave $c({\boldsymbol{x}})=A\cos(\omega_{c}{\boldsymbol{n}}\cdot{\boldsymbol{x}}+\phi_{c})$, and a message $m({\boldsymbol{x}})$ where $\hat{m}({\boldsymbol{u}})=0$ for $|{\boldsymbol{u}}|\geq\omega_{c}$, the phase modulated (PM) signal is given by:
@@
Typically the message is also sinusoidal in structure. For a baseline phase demodulation task, we attempt to recover a sinusoidal message after phase modulation and noise corruption. In this case, motivated by the phase estimation of the parabolic chirp, we use the overcomplete feature set when computing the multiscale phase to more accurately capture the overlapping frequency bands resulting from the phase modulation. After estimating the modulated phase, $\tilde{\Phi}$, it is unwrapped to produce $\tilde{\Phi}^{u}$. If successful, the message should be well estimated by either ±$\pm\tilde{\Phi}^{u}-\omega_{c}{\boldsymbol{n}}\cdot{\boldsymbol{x}}$. The $\pm$ here is due to sign ambiguity in the phase demodulation problem; for sythetic experiments the ground truth message allows us to choose the proper sign. Figure 8 shows the noisy phase modulated message, the ground truth message, and recovered messages via an amplitude based mulstiscale phase estimate versus the proposed product method. Again our tests conclude the product quality map allows for accurate phase estimation even when SNR is large.
@@ 4.3 Application to Deformable Fingerprint Registration
Global phase unwrapping algorithms typically have several major discontinuities which provide unreliable frequency information. In our experiments we compute a local frequency estimate $\omega({\boldsymbol{x}})$ by differentiation of a windowed phase unwrapping of $\Delta\phi({\boldsymbol{x}})$. The phase and orientation values used are those provided by our multiscale method. Figure 10 gives an example of the fine-scale registration produced by this method (all images from FVC2004 DB1-B [8]).
@@
Table 1 expresses the quality of the fingerprint match after successful rigid registration of fingerprints, and then after the additional fine-scale deformable registration is applied. Because we are interested only in the performance of the fine-scale registration algorithm, we add noise only after successful rigid registration. We find that even at large noise levels we are able to estimate accurate fine-scale registration in well structured areas of the fingerprints.
@@ Appendix 0.A Construction of the SMV
Two-dimensional signals can, of course, vary in two or more orientations in any given local patch, and further work has been done using hypercomplex signal processing that can deal with these cases. One extension, the structure multivector (SMV), was introduced along with the monogenic signal in the PhD dissertation of Felsberg [3]. It is designed to deal with signals of the form
@@
The features of the SMV will essentially be that of two monogenic signals, and to accommodate these additional features the SMV lives in a larger dimensional Clifford algebra, $\text{C}\ell_{3,0}$ which subsumes the quaternions. For more information on Clifford algebras see [2], [3]. In general we denote the product $e_{i_{1}}e_{i_{2}}\cdots e_{i_{n}}:=e_{i_{1}i_{2}\cdots i_{n}}$.
@@
The explicit definitions of these functions are given below, where ${\boldsymbol{x}}=xe_{1}+ye_{2}$:
$M_{1}$ $=\frac{x}{2\pi|{\boldsymbol{x}}|^{3}}*f({\boldsymbol{x}}),$ $M_{2}$ $=\frac{y}{2\pi|{\boldsymbol{x}}|^{3}}*f({\boldsymbol{x}}),$ $M_{3}$ $=f({\boldsymbol{x}}),$
$M_{23}$ $=\frac{3(3x^{2}y-y^{3})}{2\pi|{\boldsymbol{x}}|^{5}}*f({\boldsymbol{x}}),$ $M_{31}$ $=\frac{3(3xy^{2}-x^{3})}{2\pi|{\boldsymbol{x}}|^{5}}*f({\boldsymbol{x}}),$
$M_{0}$ $=\frac{-2(x^{2}-y^{2})}{2\pi|{\boldsymbol{x}}|^{4}}*f({\boldsymbol{x}}),$ $M_{12}$ $=\frac{-4xy}{2\pi|{\boldsymbol{x}}|^{4}}*f({\boldsymbol{x}}).$